OpenStudy (anonymous):

Given P(3,2,-7) and Q(4,3,-6), find the midpoint of the segment PQ. I really need help on this!!

OpenStudy (johnweldon1993):

Well it would be just like finding the midpoint of a line in 2D except here you add on the third dimension \[\large \text{Midpoint} = \frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2}, \frac{z_1 + z_2}{2}\]